# Qiuyu Wang2,    Hao Ouyang2,    Hao Chen1,    Chunhua Shen1,3,2

### v1

@@ 1 Introduction
As depicted in Fig. 1, we analyze two key metrics across four widely used benchmark datasets using two representative models: LLaDA-8B-Instruct (Nie et al., 2025) and LLaDA-1.5 (Zhu et al., 2025). The first metric, final pass rate, measures the accuracy of the final output, while the second, ever-pass rate, captures whether a correct answer appears at any point during the decoding process. In Fig. 1a, a consistent and significant discrepancy exists between these metrics. This gap reveals a critical phenomenon: models often generate correct answers during intermediate steps but subsequently overwrite them with incorrect ones. Fig. 1b illustrates this concretely—in a math problem, the model produces the correct answer “25” at sampling step 55, only to replace it with an incorrect “2” by the final step 64. More examples on temporal oscillation are presented in Sec. D.4.
@@ 2 Related Work
Diffusion language models. With the success of diffusion models in image and video generation (Song et al., 2020; Ho et al., 2020; 2022), diffusion methods have been extended to text generation. Early continuous diffusion approaches (Han et al., 2022; Li et al., 2022) operate in continuous space, while others project discrete text onto the probability simplex (Avdeyev et al., 2023; Stark et al., 2024), though these often struggle with complex linguistic patterns. More recent work applies flow matching on the simplex to learn categorical distributions (Davis et al., 2024; Cheng et al., 2024), but remains limited to simpler sequences.
@@
Discrete diffusion models, pioneered by D3PM (Austin et al., 2021), have advanced through masked token frameworks (Sahoo et al., 2024; Nie et al., 2024) and scaling efforts. Lightweight models like Plaid (Gulrajani & Hashimoto, 2023) and SEDD (Lou et al., 2023) show competitive results against GPT-2 (Radford et al., 2019), yet scalability lags behind autoregressive models. To bridge this gap, BD3-LMs (Arriola et al., 2025) and Eso-LMs (Sahoo et al., 2025) interpolate between the autoregressive and diffusion paradigms, allowing both parallel sampling and competitive performance. More recently, attempts have been made to scale up dLLMs. For example, Dream (Ye et al., 2025) converts pre-trained autoregressive models into diffusion models, while LLaDA (Nie et al., 2025) also demonstrates strong training-from-scratch results.
Another line of work explores sampling strategies in dLLMs. For example, Kim et al. (2025) examines the impact of token ordering during sampling and proposes an adaptive inference strategy that improves model performance. Similarly, ReMDM (Wang et al., 2025) introduces a remasking mechanism that leverages inference-time scaling to improve generation quality in dLLMs. Unlike prior works, we analyze and improve dLLMs from the perspective of temporal stability, a domain that remains underexplored.
@@
Test-time Strategy. Test-time strategies (Wei et al., 2022; Madaan et al., 2023; Snell et al., 2024; Yao et al., 2023; Liu et al., 2025a) for LLMs have been extensively studied and widely adopted to enhance the accuracy, consistency, and reliability of model outputs. One of the most straightforward and effective techniques in this domain is Self-Consistency (Wang et al., 2022), which selects the most consistent answer from multiple outputs using majority voting. Our approach leverages the concept of temporal self-consistency, devising a test-time strategy specifically tailored for dLLMs. Notably, our method incurs only negligible inference overhead and can be directly integrated into existing dLLMs frameworks. Another key technique is semantic entropy (Farquhar et al., 2024; Kuhn et al., 2023), a novel uncertainty metric designed for LLMs. It addresses the “semantic equivalence” problem by first clustering sentences that share the same meaning but differ in phrasing, and then computing entropy over these semantic clusters. While semantic entropy has been successfully applied to tasks like uncertainty estimation (Kuhn et al., 2023) and hallucination detection (Farquhar et al., 2024), we introduce Temporal Semantic Entropy to characterize the stability and confidence of dLLMs during the progressive denoising process.
@@
Post-Training using Reinforcement Learning. Group Relative Policy Optimization (Shao et al., 2024; Guo et al., 2025) (GRPO) is a variant of Proximal Policy Optimization (Schulman et al., 2017) (PPO) that directly computes the advantage from rewards within a group of responses, obviating the need for an separately trained value function. GRPO has demonstrated not only strong performance on reasoning tasks such as mathematics and code generation, but also significant potential across a broader range of modalities and downstream tasks (Huang et al., 2025; Shen et al., 2025; Qi et al., 2025; Zhong et al., 2025; Li et al., 2025; Damani et al., 2025). Building on this, related works have introduced further refinements. For instance, DAPO (Yu et al., 2025) employs a dynamic sampling strategy to prevent training batches from consisting entirely of correct or incorrect samples. Other approaches (Zhang et al., 2025; Prabhudesai et al., 2025; Cui et al., 2025; Agarwal et al., 2025) have enhanced RL strategies from an entropy perspective; EMPO (Zhang et al., 2025) estimates rewards directly based on semantic entropy, while Seed-GRPO (Chen et al., 2025) leverages entropy to improve advantage estimation. Recently, the GRPO framework has been adapted for dLLMs in methods like diffu-GRPO (Zhao et al., 2025), UniGRPO (Yang et al., 2025), and coupled-GRPO (Gong et al., 2025). A commonality among these methods is their reliance on ground-truth answers to formulate reward signals. In contrast, our approach is distinguished by its unsupervised nature; it enhances the model’s temporal consistency without depending on ground-truth supervision.
@@ 3.1 Preliminaries on dLLMs
This generative process can be divided into two key steps. The first step involves $p_{\theta}(x_{0}\mid x_{t})$, a model trained to predict the clean text sequence based on the current noisy input. We denote the answer decoded at the $t$-th step as $x_{0}^{t}$, i.e., $x_{0}^{t}=p_{\theta}(x_{0}\mid x_{t})$. The second step relies on $q(x_{t-1}\mid x_{0})$, referred to as the re-masking process. This process applies the forward procedure to the currently predicted $x_{0}^{t}$, yielding a sequence $x_{t-1}$ that is less noisy compared to $x_{t}$ at time $t$. Various strategies can be employed for this process, such as random re-masking or low-confidence re-masking, as discussed in (Nie et al., 2025).
@@ 3.2 Temporal Oscillation
In the reverse generative process of dLLMs, the prediction of the clean sequence $p_{\theta}(x_{0}\mid x_{t})$ at a single denoising step is often inaccurate, especially when the noise level $t$ is large. Existing models perform iterative denoising according to the generative process framework, where the final output is determined by the prediction $p_{\theta}(x_{0}\mid x_{1})$ at the last denoising step, while neglecting all intermediate predictions $\{x_{0}^{t}=p_{\theta}(x_{0}\mid x_{t})\}_{t=2}^{T}$ generated during the iterative process. In this work, we conduct an in-depth investigation into these intermediate step results, revealing a critical phenomenon in diffusion-based text generation.
@@
Experiment Setup. To validate the existence of fluctuating correctness across timesteps, we evaluate both the final pass rate, i.e., the model’s $\operatorname{Pass}@1$ at the last denoising step, and the $\operatorname{EverPass}@1\mid t$ for two representative dLLMs: LLaDA-8B-Instruct (Nie et al., 2025) and LLaDA-1.5 (Zhu et al., 2025). The evaluation is conducted on varying answer length and four representative mathematical reasoning benchmarks: GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023), SVAMP (Patel et al., 2021), and Countdown (Pan et al., 2025).
Figure 2: Patterns of accuracy evolution over diffusion sampling steps. Responses are generated with a length of 128 tokens using 64 diffusion steps, based on the LLaDA-8B-Instruct model. Left: Accuracy generally improves with more sampling steps across all datasets. Simpler datasets like SVAMP start high, while harder ones like Countdown begin lower but steadily improve. Middle and Right: We compare the final pass rate, $\operatorname{Pass}@1$, with the cumulative “ever pass rate” over steps, $\operatorname{EverPass}@1\mid t$. A consistent gap remains between $\operatorname{Pass}@1$ and $\operatorname{EverPass}@1\mid t$, highlighted by the green shaded area in the figure.
Observations. As shown in Fig. 1 and Table 1 in Sec. 5.2, there exists a significant gap between the final pass rate and the ever pass rate . For example, on the GSM8K dataset with a sequence length of 128, the final pass rate of LLaDA-8B-Instruct is 68.5%, while its ever pass rate reaches 80.5%, resulting in a gap of 12.0%. This discrepancy indicates that a considerable number of questions are correctly answered in one or more intermediate steps, but the answers are subsequently modified to be incorrect during later stages of the refinement process. This phenomenon highlights a potential instability in the model’s iterative decoding, where correct intermediate paths may be overwritten or lost as generation progresses. We term this phenomenon temporal oscillation, with examples presented in Sec. D.4.
@@ 3.3 Analyses
To gain a deeper understanding of the temporal oscillation phenomenon, we conduct comprehensive analyses from multiple dimensions: the evolution of accuracy, the dynamics of entropy, and the semantic stability across the decoding process.
Accuracy Across Sampling Steps. To better understand the temporal oscillation phenomenon, we first visualize how accuracy evolves across diffusion sampling steps. As shown in Fig. 2a, overall, accuracy improves with increasing sampling steps across all datasets. Furthermore, the models exhibit distinct accuracy patterns depending on task complexity. For simpler datasets like SVAMP, accuracy is already relatively high at the initial steps. Correct answers often appear early and are gradually refined, with minor fluctuations. In contrast, for more complex datasets like Countdown, the models tend to struggle initially, i.e., accuracy starts low, but they benefit significantly from iterative refinement, with performance improving steadily over subsequent steps.
To better understand these dynamics, we further analyze GSM8K and Countdown, computed $\operatorname{Pass}@1$ and $\operatorname{EverPass}@1\mid t$ for all intermediate steps. While initial correct predictions appear early in GSM8K but later in Countdown, a noticeable gap between $\operatorname{Pass}@1$ and $\operatorname{EverPass}@1\mid t$ emerges and gradually widens as the sampling process unfolds. These results suggest that early correct predictions do not guarantee stable reasoning, highlighting the importance of preserving correct intermediate states throughout the sampling process. More analysis on the SVAMP and MATH500 datasets is presented in Sec. C.1.
@@
Figure 3: Patterns of entropy evolution over diffusion sampling steps. Responses are generated with a length of 128 using 64 diffusion steps, based on the LLaDA-8B-Instruct model. Left: The average token-level entropy decreases steadily during the sampling process. GSM8K exhibits lower entropy than Countdown, consistent with the model’s better performance on GSM8K. Middle and Right: Both Intermediate-Correct and Always-Incorrect questions exhibit higher overall entropy compared to Finally-Correct ones. On GSM8K, Intermediate-Correct questions display lower entropy in the early steps than Always-Incorrect, indicating initial confidence, whereas on Countdown the entropy trend is less stable. Note that some points in Right are missing due to the inability to extract answers from generated responses at early diffusion steps for Countdown.
Entropy Across Sampling Steps. Temporal oscillations reflect the model’s uncertainty during generation. To quantify this, we analyze the average token-level entropy across sampling steps. According to Fig. 3, we observe that across all datasets, the average token-level entropy decreases steadily during the sampling process and approaches zero at the final step. Notably, GSM8K exhibits relatively low entropy from the beginning, indicating higher initial confidence in its predictions.
To understand how entropy relates to correctness, we further categorize questions into three groups: (1) Finally-Correct: the final answer is correct, regardless of intermediate steps; (2) Intermediate-Correct: at least one intermediate step is correct, but the final answer is incorrect; (3) Always-Incorrect: all sampling steps, including the final answer, are incorrect. We compute the average entropy of tokens associated with the predicted answers across diffusion steps for GSM8K and Countdown. The results in Fig. 3a show that questions with incorrect final answers, covering both the Always-Incorrect and Intermediate-Correct categories, tend to exhibit higher overall entropy. Notably, GSM8K questions in the Intermediate-Correct category show lower entropy than Always-Incorrect ones in the early steps, as shown in Fig. 3b, suggesting initial prediction that deteriorates during the refinement process.
Considering the use of a semi-autoregressive sampling strategy (Nie et al., 2025), we additionally measure the average entropy of the currently generated block, detailed in Sec. C.3.
@@
Figure 4: Temporal semantic entropy across four mathematical benchmarks. This metric measures the uncertainty in the semantic content of answers throughout the decoding steps. Statistically, correctly answered questions exhibit lower entropy.
@@
Temporal Semantic Entropy. While token-level entropy measures local uncertainty, we need a metric that captures semantic consistency across the entire decoding trajectory. To this end, we introduce a metric called Temporal Semantic Entropy. It captures the distribution of semantic variations in answers generated at each step of the iterative denoising process.
@@
During decoding, we obtain a sequence of $T$ intermediate answers, denoted by $\{x_{0}^{t}\}_{t=1}^{T}$. We group these answers into clusters based on their semantic meaning, forming a set of semantic clusters $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{K}\}$, where each cluster $C_{k}=\{x_{0}^{t}:\text{meaning}(x_{0}^{t})=k\}$ contains all answers with equivalent semantics (outcome equals $k$). We then define the temporal semantic entropy (TSE) of a trajectory as:
$\operatorname{TSE}(\left\{x_{0}^{t}\right\}_{t=1}^{T})=-\sum_{C_{k}}\left(\left[\sum_{x_{0}^{t}\in C_{k}}p\left(x_{0}^{t}\right)\right]\log\left[\sum_{x_{0}^{t}\in C_{k}}p\left(x_{0}^{t}\right)\right]\right),$
@@
which quantifies the uncertainty in the semantic content of answers over the decoding steps. A higher value indicates more semantic variation throughout the trajectory, while a lower value suggests convergence toward a consistent meaning.
@@
As illustrated in Fig. 4, temporal semantic entropy provides valuable insight into the model’s behavior during the generation process. In datasets like Countdown and MATH, where model performance is comparatively lower, we observe a higher temporal semantic entropy than in GSM8K and SVAMP, suggesting increased semantic instability and uncertainty throughout the generation trajectory. Moreover, questions that are ultimately answered correctly, encompassing both Intermediate-Correct and Always-Incorrect, tend to exhibit lower temporal semantic entropy than those answered incorrectly. This indicates that more stable and semantically consistent decoding trajectories are associated with better performance. Consequently, high temporal semantic entropy may serve as a signal for model uncertainty and could be used to prioritize samples for further training or refinement. We refer readers to Sec. C.2 for more observations on temporal semantic entropy.

### v2

@@ 1 Introduction
As depicted in Fig. 1, we analyze two key metrics across four widely used benchmark datasets using two representative models: LLaDA-8B-Instruct (Nie et al., 2025) and LLaDA-1.5 (Zhu et al., 2025). The first metric, final pass rate, measures the accuracy of the final output, while the second, ever-pass rate, captures whether a correct answer appears at any point during the decoding process. In Fig. 1a, a consistent and significant discrepancy exists between these metrics. This gap reveals a critical phenomenon: models often generate correct answers during intermediate steps but subsequently overwrite them with incorrect ones. Fig. 1b illustrates this concretely—in a math problem, the model produces the correct answer “25” at sampling step 55, only to replace it with an incorrect “2” by the final step 64. More examples on temporal oscillation are presented in Sec. F.4.
@@ 2 Related Work
Diffusion language models. Building on the success of diffusion in image and video generation (Song et al., 2020; Ho et al., 2020; 2022), diffusion methods have been extended to text. Early continuous approaches (Han et al., 2022; Li et al., 2022) operate in continuous space, while others map text to the probability simplex (Avdeyev et al., 2023; Stark et al., 2024). More recent work applies flow matching on the simplex to learn categorical distributions (Davis et al., 2024; Cheng et al., 2024), but remains limited to simpler sequences.
@@
Discrete diffusion models, pioneered by D3PM (Austin et al., 2021), advanced through masked token frameworks (Sahoo et al., 2024; Nie et al., 2024) and scaling efforts. Lightweight variants like Plaid (Gulrajani & Hashimoto, 2023) and SEDD (Lou et al., 2023) rival GPT-2 (Radford et al., 2019), yet lag autoregressive models in scalability. To bridge this gap, BD3-LMs (Arriola et al., 2025) and Eso-LMs (Sahoo et al., 2025) interpolate between autoregressive and diffusion paradigms, enabling parallel sampling with competitive performance. Recent efforts scale up dLLMs: Dream (Ye et al., 2025) converts pretrained autoregressive models into diffusion models, while LLaDA (Nie et al., 2025) trains strong models from scratch.
Another line of work studies sampling in dLLMs. For instance, Kim et al. (2025) show token ordering affects performance and propose adaptive inference, while ReMDM (Wang et al., 2025) uses inference-time remasking to boost generation. Unlike previous work, we focus on improving dLLMs from the underexplored perspective of temporal stability.
@@
Test-time Strategy. Test-time strategies (Wei et al., 2022; Madaan et al., 2023; Snell et al., 2024; Yao et al., 2023; Liu et al., 2025a) are widely used to improve LLM accuracy, consistency, and reliability. A simple yet effective method is Self-Consistency (Wang et al., 2022), which selects the most consistent answer from multiple outputs via majority voting. Building on this idea, we propose a temporal self-consistency strategy tailored for dLLMs, which adds negligible inference overhead and integrates seamlessly into existing frameworks. Another important technique is semantic entropy (Farquhar et al., 2024; Kuhn et al., 2023), an uncertainty metric that clusters semantically equivalent outputs before computing entropy. While previously applied to uncertainty estimation and hallucination detection, we extend it to dLLMs by introducing Temporal Semantic Entropy, capturing stability and confidence throughout the denoising process.
@@
Post-Training using Reinforcement Learning. Group Relative Policy Optimization (Shao et al., 2024; Guo et al., 2025) (GRPO), a variant of Proximal Policy Optimization (Schulman et al., 2017), computes advantages directly from group rewards, removing the need for a separately trained value function. GRPO has shown strong performance in reasoning tasks like mathematics and code generation, and promising results across broader modalities (Huang et al., 2025; Shen et al., 2025; Qi et al., 2025; Zhong et al., 2025; Li et al., 2025; Damani et al., 2025). Building on this, refinements such as DAPO (Yu et al., 2025) introduce dynamic sampling to balance training batches, while entropy-based methods (Zhang et al., 2025; Prabhudesai et al., 2025; Cui et al., 2025; Agarwal et al., 2025) further enhance RL. For example, EMPO (Zhang et al., 2025) derives rewards from semantic entropy, and Seed-GRPO (Chen et al., 2025) improves advantage estimation. Recent adaptations to dLLMs include diffu-GRPO (Zhao et al., 2025), UniGRPO (Yang et al., 2025), and coupled-GRPO (Gong et al., 2025), which still rely on ground-truth rewards. By contrast, our method is fully unsupervised, enhancing temporal consistency without ground-truth supervision.
@@ 3.1 Preliminaries on dLLMs
This generative process can be divided into two key steps. The first step involves $p_{\theta}(x_{0}\mid x_{t})$, a model trained to predict the clean text sequence based on the current noisy input. We denote the answer decoded at the $t$-th step as $x_{0}^{t}$, i.e., $x_{0}^{t}=p_{\theta}(x_{0}\mid x_{t})$. The second step relies on $q(x_{t-1}\mid x_{0})$, referred to as the re-masking process. This process applies the forward procedure to the currently predicted $x_{0}^{t}$, yielding a sequence $x_{t-1}$ that is less noisy compared to $x_{t}$. Various strategies can be employed for this process, such as random or low-confidence re-masking, as in (Nie et al., 2025).
@@ 3.2 Temporal Oscillation
In the reverse process of dLLMs, predictions $p_{\theta}(x_{0}\mid x_{t})$ at a single step are often inaccurate, especially under high noise when $t$ is large. Existing models perform iterative denoising according to the generative process framework, where the final output is determined by the prediction $p_{\theta}(x_{0}\mid x_{1})$ at the last denoising step, while neglecting all intermediate predictions $\{x_{0}^{t}=p_{\theta}(x_{0}\mid x_{t})\}_{t=2}^{T}$ generated during the iterative process. In this work, we conduct an in-depth investigation into these intermediate step results, revealing a critical phenomenon in diffusion-based text generation.
@@
Experiment Setup. We compare the final pass rate, i.e., $\operatorname{Pass}@1$ at the last step with $\operatorname{EverPass}@1\mid t$ on two representative dLLMs: LLaDA-8B-Instruct and LLaDA-1.5, evaluated across different answer lengths and four reasoning benchmarks: GSM8K (Cobbe et al., 2021), MATH500 (Lightman et al., 2023), SVAMP (Patel et al., 2021), and Countdown (Pan et al., 2025).
Figure 2: Patterns of accuracy evolution over diffusion sampling steps. Responses of length 128 are generated with 64 steps using LLaDA-8B-Instruct. Left: Accuracy generally rises with more steps across datasets; SVAMP starts high, while harder ones like Countdown start low but improve steadily. Middle/Right: We compare the final pass rate, $\operatorname{Pass}@1$, with cumulative $\operatorname{EverPass}@1\mid t$ over steps. A clear gap persists between them, shown by the green shaded area.
Observations. As shown in Fig. 1 and Table 1, there is a notable gap between the final pass rate and the ever pass rate . For instance, on GSM8K with length 128, LLaDA-8B-Instruct achieves 68.5% final pass rate versus 80.5% ever pass rate , a gap of 12.0%. This gap shows that many questions are correctly solved at intermediate steps but later revised to incorrect answers during refinement. It reveals an instability in iterative decoding, where correct paths can be overwritten as generation proceeds. We term this phenomenon temporal oscillation, with examples in Sec. F.4.
@@ 3.3 Analyses
To gain a deeper understanding of the temporal oscillation phenomenon, we conduct comprehensive analyses from multiple dimensions: accuracy, entropy, and semantic stability across decoding steps.
Accuracy Across Sampling Steps. As shown in Fig. 2a, accuracy generally improves with more sampling steps. Simpler datasets like SVAMP start high and remain stable, while harder ones like Countdown begin low but benefit from iterative refinement. To probe further, we compare $\operatorname{Pass}@1$ and $\operatorname{EverPass}@1\mid t$ on GSM8K and Countdown. Early correct predictions appear sooner on GSM8K but later on Countdown, and a growing gap between the two metrics reveals that early correctness does not ensure stable reasoning. This underscores the importance of preserving correct intermediate states. Additional results for SVAMP and MATH500 are in Sec. D.1.
@@
Figure 3: Patterns of entropy evolution over diffusion sampling steps. Responses are generated with length 128 using 64 diffusion steps from the LLaDA-8B-Instruct model. Left: Average token-level entropy decreases steadily during sampling. GSM8K shows lower entropy than Countdown, aligning with its higher accuracy. Middle and Right: Both Intermediate-Correct and Always-Incorrect questions exhibit higher overall entropy compared to Finally-Correct ones. On GSM8K, Intermediate-Correct questions display lower entropy in the early steps than Always-Incorrect, indicating initial confidence, whereas on Countdown the entropy trend is less stable.
@@
Figure 4: Temporal semantic entropy across four benchmarks. This metric measures the uncertainty in the semantic content of answers across decoding steps. Statistically, correctly answered questions exhibit lower entropy.
Entropy Across Sampling Steps. Temporal oscillations reflect model uncertainty. To quantify this, we analyze the average token-level entropy. As shown in Fig. 3, entropy decreases steadily and approaches zero by the final step, with GSM8K starting lower, indicating higher initial confidence. We categorize questions into three groups: Finally-Correct, Always-Incorrect, and Intermediate-Correct, where Intermediate-Correct means at least one intermediate step is correct but the final answer is incorrect. For GSM8K and Countdown, incorrect final answers show consistently higher entropy. On GSM8K, Intermediate-Correct questions begin with lower entropy than Always-Incorrect, suggesting initially confident but unstable predictions. Considering the semi-autoregressive sampling strategy (Nie et al., 2025), we also measure the average entropy of the currently generated block, detailed in Sec. D.3.
@@
Temporal Semantic Entropy. Token-level entropy reflects local uncertainty, but we also need a measure of semantic consistency across the decoding trajectory. We therefore introduce Temporal Semantic Entropy (TSE), which captures the semantic variations of answers during sampling.
@@
During decoding, we obtain a sequence of $T$ intermediate answers, denoted by $\{x_{0}^{t}\}_{t=1}^{T}$. We cluster them by semantic meaning into $\mathcal{C}=\{C_{1},\ldots,C_{K}\}$, where each cluster $C_{k}$ groups answers with equivalent semantics. The TSE of a trajectory is defined as:
$\operatorname{TSE}(\{x_{0}^{t}\}_{t=1}^{T})=-\sum_{C_{k}}\Big(\!\!\sum_{x_{0}^{t}\in C_{k}}p(x_{0}^{t})\Big)\log\Big(\!\!\sum_{x_{0}^{t}\in C_{k}}p(x_{0}^{t})\Big),$
@@
which quantifies the uncertainty in semantic content across steps: higher TSE indicates more semantic variation, while lower TSE implies convergence to a consistent meaning.
@@
As shown in Fig. 4, TSE offers insight into model behavior during generation. In datasets like Countdown and MATH, where performance is weaker, we observe higher entropy than in GSM8K and SVAMP, reflecting greater semantic instability. Moreover, questions ultimately answered correctly generally exhibit lower entropy than incorrect ones, indicating that stable, semantically consistent trajectories align with better performance. Thus, high TSE may signal model uncertainty and highlight samples for further improvement. More results on TSE are provided in Sec. D.2.